Find the inverse Laplace Transform of {1/s^3(s^2+1)}


1
Expert's answer
2022-01-14T08:05:39-0500

We shall find the inverse laplace transform analytically of

"\displaystyle\frac{1}{s^{3}(s^{2}+1)}"

We shall first resolve the expression by partial fraction

"\displaystyle\frac{1}{s^{3}(s^{2}+1)}=\frac{A}{s}+\frac{B}{s^{2}}+\frac{C}{s^{3}}+\frac{Ds+E}{s^{2}+1}"


"1\equiv As^{2}(s^{2}+1)+Bs(s^{2}+1)+C(s^{2}+1)+Ds^{4}+Es^{3}"

It follows that

"[s^{4}]" "A+D=0"

"[s^{3}]" "B+E=0"

"[s^{2}]" "A+C=0"

"[s]" "B=0 \implies E=0"


We have that "C=1\implies A=-1, D=1"


NOW, we have

"\displaystyle\frac{1}{s^{3}(s^{2}+1)}=\frac{-1}{s}+\frac{1}{s^{3}}+\frac{s}{s^{2}+1}"


"\mathcal{L}^{-1}\left\{\displaystyle\frac{1}{s^{3}(s^{2}+1}\right\}=\mathcal{L}^{-1}\left\{\frac{-1}{s}+\frac{1}{s^{3}}+\frac{s}{s^{2}+1}\right\}"


"=\mathcal{L}^{-1}\left\{\frac{-1}{s}\right\}+\mathcal{L}^{-1}\left\{\frac{1}{s^{3}}\right\}+\mathcal{L}^{-1}\left\{\frac{s}{s^{2}+1}\right\}"


"=-1+\frac{t^{2}}{2}+cos(t)"

"=cos(t)+\frac{t^{2}}{2}-1"

which is the inverse laplace transform of "\displaystyle\frac{1}{s^{3}(s^{2}+1)}"